OpenStudy (anonymous):

A system of equations is shown below: -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

I only need help with part B

OpenStudy (campbell_st):

ok... so what solution did you get..?

OpenStudy (anonymous):

for part A?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

yes can you post the solution to part A

OpenStudy (anonymous):

-6x+14y=-32 I got this I multiplied first equation by 2

OpenStudy (campbell_st):

ok.... this is what I would have done I would multiply every term in the 1st equation by -3. this would allow you to eliminate x equation 2 doesn't change 9x - 21y = 48 -9x +5y = 16 add the 2 equations --------------- -16y = 64 so y = -4 substitute this into either of the original equations to find x. does that make sense...?

OpenStudy (anonymous):

yes thanks you soo much

OpenStudy (campbell_st):

ok... so when you get the solution part B seems to be asking you to substitute the values into 9x - 21y = 48 to show the solution is correct